Problem 7-8B Record the disposal of equipment (LO7-6)

[The following information applies to the questions displayed below.]

New Deli is in the process of closing its operations. It sold its three-year-old ovens to Sicily Pizza for $320,500. The ovens originally cost $427,500, had an estimated service life of 10 years, had an estimated residual value of $27,500, and were depreciated using straight-line depreciation. Complete the requirements below for New Deli.

Required information Problem 7-8B Record the disposal of equipment (LO7-6) [The following information applies to the questionNew Deli is in the process of closing its operations. It sold its three-year-old ovens to Sicily Pizza for $320,500. The ovenProblem 7-8B Part 3 3. What is the gain or loss on the sale of the ovens at the end of the third year?4. Record the sale of the ovens at the end of the third year. (If no entry is required for a transaction/event, select No JoJournal entry worksheet 1 Record the disposal of equipment. Note: Enter debits before credits. Transaction General Journal De
Answer #1

Depreciation = (Cost - Salvage value) /Useful life

1) Accumulated Depreciation

= (427,500-27,500)/10 * 3

= 120,000

​​​​​​2) Book value = Cost - Accumulated Depreciation

= 427,500 - 120,000

= 307,500

3) Gain (Loss) = Cash received - Book value

= 320,500 - 307,500

Gain of 13,000

Cash 320,500
Accumulated Depreciation - oven 120,000
Gain on sale of oven 13,000
Oven 427,500